[Koji / Kiwamu]
The Michelson was locked with the new LSC realtime code.
(what we did)
-- Fine alignment of the Michelson, including PZTs, BS and ITMY.
Since the X arm has been nicely aligned we intentionally avoided touching ITMX. The IR beam now is hitting the center of both end mirrors.
At the end we lost X arm's resonance for IR. This probably means the PZTs need more careful alignments.
-- Signal acquisition
We replaced the RFPD (AS55) that Aidan and Jamie nicely installed by POY11 because we haven't yet installed a 55MHz RF source.
The maximum DC voltage from the PD went to about 50 mV after aligning steering mirrors on the AP table.
The RF signal from the PD is transferred by a heliax cable which has been labeled 'REFL33'.
Then the RF signal is demodulated at a demodulation board 'AS11', which is one of the demodulation boards that Suresh recently modified.
Although we haven't fully characterized the demod board the I and Q signal looked healthy.
Finally the demod signals go to ADC_0_3 and ADC_0_4 which are the third and fourth channel.
They finally show up in REFL33 path in the digital world.
With the new LSC code we fedback the signal to BS. We put anti-whitening filters in the I and Q input filter banks.
We found that dataviewer didn't show correct channels, for example C1LSC_NREFL33I showed just ADC noise and C1LSC_NREFL33Q showed NREFL_33I.
Due to this fact we gave up adjusting the digital phase rotation and decided to use only the I-phase signal.
Applying a 1000:10 filter gave us a moderate lock of the Michelson. The gain was -100 in C1LSC_MICH_GAIN and this gave us the UGF of about 300 Hz.
Note that during the locking both ETMs were intentionally misaligned in order not to have Fabry-Perot fringes.